∫f(×)+ ∫g(×) . What is the value of the integral is the integration is running from 0 to a for f(×) . And the integration is running from 0 to b. For g(×).

∫f(×)+  ∫g(×) .          What is the value of the integral is the integration is running from 0 to a for f(×) .        And the integration is running from 0 to b.         
For g(×).
 

Grade:10

1 Answers

Arun
5 years ago
Dear Trishna
 
∫f(×) = F (x)
then definite integral = F (b) – F(a)
and 
if ∫g(×) = G (x)
hence
G(b) – G(a)

Regards
Arun